# Ctrl-Crash: Controllable Diffusion for Realistic Car Crashes

### v1

@@ Abstract
Figure 1: Counterfactual Crash Generation: this diagram demonstrates the ability of our model to generate counterfactual crashes (middle-row: no crash, bottom-row: ego/car crash) while beginning from the initial frame and 3 bounding-boxes frames of the real video (top-row: the real car crash).
@@ 1 Introduction
Autonomous vehicle (AV) systems must be rigorously tested in a wide range of driving situations—including rare and dangerous edge cases such as collisions to ensure safe deployment. Much of the current progress in perception, planning, and control for AVs has been driven by large-scale datasets collected in harmless, non-crash scenarios. However, realistic video data of car crashes remains extremely scarce, making it difficult to simulate, anticipate, or learn effectively from these critical events (wang2023deepaccident, ; sadegh2018viena2, ; kim2019aaai, ).
Prior work (luo2025ctrlv, ; wang2024drivedreamer, ; zhao2025drivedreamer, ; kim2021drivegan, ; hu2023gaia1, ) has largely approached the challenge of crash scenario modeling in two main ways. On one hand, physics-based rendering approaches use game engines or physics simulators to model accident dynamics, but often fall short on visual realism, require expensive rendering pipelines, and demand costly human effort for environment and asset creation (dosovitskiy17carla, ; cabon2020virtualkitti2, ; lehner2022_3d-vfield, ). On the other hand, data-driven methods, such as generative models, rely on real-world footage, which is difficult to acquire in sufficient volume due to the infrequency and ethical complexity of crash events (Geiger2013kitti, ; dosovitskiy17carla, ; holger2019nuscnes, ). Moreover, most generative approaches focus on normal driving behavior, avoiding the complexity and unpredictability inherent in crash dynamics, where agent interactions, rare motion patterns, and semantic context all matter deeply (wu2024drivescape, ; zhao2025drivedreamer, ; luo2025ctrlv, ; kim2021drivegan, ).
@@
Ctrl-Crash builds on latent diffusion models (rombach2022high, ) and classifier-free guidance (ho2022cfg, ), and we extend the latter to allow independently tunable guidance strengths for each control modality, making our system highly flexible at inference time. Our two-stage training procedure first finetunes a pretrained Stable Video Diffusion (SVD) (blattmann2023svd, ) model on in-the-wild ego-view accident videos, then trains a ControlNet (zhang2023controlnet, ) adapter to handle conditioning in order to direct the video generation. By leveraging this data-driven framework, our model can generate controllable crash videos that are visually realistic, semantically diverse, and behaviorally plausible. We see this work as a step toward not only improving the diversity and coverage of safety-critical AV testing, but also enabling counterfactual safety reasoning: the ability to simulate alternate outcomes from identical initial conditions, and better understand the causality of crashes.
@@
    We develop a data-processing pipeline to filter driving videos and extract bounding box trajectories of road users. Using it, we release a curated extension of the MM-AU dataset (fang2024abductive, ) (with bounding box annotations), as well as held-out test sets from Car Crash Dataset RUSSIA (sivoha2023carcrash, ) (with computed boxes) and BDD100k (yu2020bdd100k, ) (with existing boxes), along with tools to support research in crash simulation and controllable video generation.
@@ 2 Related Work
Video Diffusion Models. Diffusion models (song2019generative, ; sohl2015diffusion, ; ho2020ddpm, ) have emerged as a powerful paradigm for generative modeling, particularly in the domain of image synthesis. They operate by learning to reverse a gradual noising process applied to data, generating high-fidelity samples through iterative denoising. Recent advances have extended these techniques to the video domain, where temporal consistency and spatial coherence are critical (ho2022vdm, ; he2023lvdm, ; blattmann2023svd, ).
Given the noise-level $t$, and task-specific conditions $\bm{c}$, the diffusion loss is the following:
$\vspace{-3pt}\mathcal{L}=\mathbb{E}_{\bm{z}_{0},\bm{t},\bm{c}_{t},\bm{c}_{\text{f}},\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(x_{t},t,c))\|_{2}^{2}\Big{]},$ (1)
where $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, $\epsilon\sim\mathcal{N}(0,1)$, and $\bar{\alpha}_{t}$ at $t\in[1,T]$ controls the diffusion schedule.
Latent video diffusion models (LVDMs) (he2023lvdm, ; blattmann2023svd, ) address the computational challenges of high-resolution video generation by operating in a compressed latent space. This enables the generation of long, high-quality video sequences from compact representations. Specifically, stable video diffusion (SVD) (blattmann2023svd, ), a variant of LVDM, leverages a UNet-based denoiser trained on video latents conditioned on an initial frame, making it suitable for tasks like image-to-video generation and temporal extension.
@@
Controllable Generative Models. Recent progress in generative modeling has emphasized not only fidelity, but also controllability — the ability to guide outputs through structured input signals. In image generation, this includes control via text prompts, sketches, bounding boxes, semantic maps, and more. In diffusion models, Classifier-Free Guidance (CFG) (ho2022cfg, ) and ControlNet (zhang2023controlnet, ) have been introduced as effective methods to allow adaptable conditioning while preserving high-quality generation.
Classifier-Free Guidance is a widely-used technique in diffusion models to improve conditional generation by combining conditional and unconditional predictions, scaled to control how strongly the model follows the conditioning input. It involves jointly training the diffusion model for conditional and unconditional denoising by randomly setting the conditioning to a null value $c=\emptyset$ during training. During inference, the denoising prediction is interpolated between a conditional noise estimate $\epsilon_{\theta}(\mathbf{x}_{t},c)$ and an unconditional noise estimate $\epsilon_{\theta}(\mathbf{x}_{t},\emptyset)$, scaled by a guidance factor $\gamma$ to obtain the modified score estimate:
@@
Vision models, such as InstructPix2Pix (pix2pix, ), use textual or mixed-mode conditioning with CFG to manipulate generation intent and content mid-sampling. ControlNet (zhang2023controlnet, ) introduced an effective approach for injecting spatial control into diffusion models by adding a parallel, trainable network that processes conditioning signals and modulates the main denoising backbone. It works by branching off from intermediate layers in the main U-Net and processing a control input in parallel. Its outputs are then added to the original U-Net features before the denoising step, effectively steering the generation without retraining the base model
@@
Car Crash Simulation and AV Safety. Video diffusion models offer a compelling solution for car crash simulation, as they can simulate both the visual realism and behavioral dynamics of complex driving scenes with very little user effort. Recent car video generation models (wang2024drivedreamer, ; zhao2025drivedreamer, ; luo2025ctrlv, ; agarwal2025cosmos, ) focus on structured driving video generation using control signals like bounding boxes or text prompts. These models generate temporally consistent driving scenes and support the synthesis of structured traffic interactions. However, they are typically limited to non-crash scenarios or coarse control.
Physics-based crash simulations (dosovitskiy17carla, ; gabbard2019ar, ) are crucial for evaluating autonomous vehicle safety (dupuy2001generating, ; chen2021geosim, ; rempe2022generating, ; wang2024deepaccident, ). Due to the scarcity and ethical challenges of real-world crash data, there is growing interest in generating synthetic safety-critical scenarios (wang2024deepaccident, ; huang2025safety, ). These simulations provide control and physical realism but often lack authenticity and broad applicability. In contrast, our work focuses on rare, adversarial cases that challenge perception and planning systems in real-world scenarios. To address the data challenge, we propose a novel method of data processing.
CtRL-Sim rowe2024ctrlsim uses return conditioned offline reinforcement learning to model AV agent control from real-world trajectories. By applying a negative exponential tilt, it simulates realistic crash dynamics as agents pursue collision-related negative rewards. Operating at the data-structure level, it predicts 2D bounding box trajectories in bird’s eye views using simplified “bicycle models." While useful for generating synthetic crash data, real-world crash dynamics are more complex than those produced by this approach.
While most prior work has focus on general driving scenes, notable examples of prior work tackling realistic car crash generation using real videos include: DrivingGen (guo2024drivinggen, ), which generates crash videos from textual accident reports, AVD2 (li2025avd2, ) and OAVD (fang2024abductive, ) which also generate videos of car crashes, but their main focus is car crash video description rather than generation. These methods are designed to stress-test AV policies and enhance safety coverage beyond what real-world datasets offer.
Our work builds on this vision by proposing a controllable video diffusion model that can generate crash outcomes directly from initial conditions and desired outcomes. Unlike physics simulators, Ctrl-Crash supports semantic control (e.g., crash type) and spatial trajectory specification (e.g., bounding boxes), enabling targeted stress testing of vision-based AV stacks. Moreover, our method supports counterfactual safety reasoning, allowing one to explore how small changes in agent behavior or intent could lead to dramatically different outcomes—a critical capability for understanding near-miss events and decision boundary failures. Importantly, our method has much higher visual and motion fidelity than prior works (See Section 4).
@@ 3 Our Method: Ctrl-Crash
Figure 2: Model architecture: Ctrl-Crash treats Bounding Boxes (BBs) as images. Both BBs and images frames as fed to a VAE image encoder. The crash type and BB embeddings are fed to ControlNet. The images embedding after adding noise ($x_{t}$), the noise level ($t$), and the ControlNet intermediate outputs ($c$) are fed to the Stable Video Diffusion (SVD) model to obtain the predicted noise $\epsilon_{\theta}(x_{t},t,c)$. Clip embeddings are computed by passing the first image for each video through a pretrained CLIP encoder. These CLIP embeddings are then fed to the ControlNet and SVD models. on The diffusion process is solved over multiple steps using Classifier-Free Guidance in the latent space and then decoded back to images using the VAE image decoder. See Section 3 for details.
@@ 3.1 Overview
We propose Ctrl-Crash (Figure 2), a controllable video diffusion framework designed to generate realistic car crash scenarios from a single initial frame, guided by both spatial and semantic control signals. Ctrl-Crash builds on Ctrl-V luo2025ctrlv , a framework for generating videos from rendered bounding box trajectories, by extending its capabilities to crash-specific scenarios, offering richer control and greater flexibility. Specifically, we incorporate a new semantic control signal representing crash type and introduce a refined training procedure to handle partial and noisy conditioning.
Our method follows a two-stage training pipeline. In the first stage, we fine-tune a Stable Video Diffusion (SVD) model on the MM-AU (fang2024abductive, ) crash video dataset to improve its ability to synthesize dynamic and physically plausible driving scenes. In the second stage, we train a ControlNet module to inject conditioning information in two forms: (1) bounding box sequences representing road user motion across time, and (2) discrete crash type labels encoding high-level semantic intent. To ensure generalization to incomplete or noisy control, we introduce a curriculum-based random masking strategy that progressively masks out parts of the control inputs during training. Masked bounding box frames are replaced by a learnable null embedding that preserves scene plausibility. We further extend classifier-free guidance to support independent scaling of each control modality, enabling nuanced and flexible control at inference. Unconditional noise predictions are obtained from the pretrained base model for improved generation diversity and stability.
Ctrl-Crash supports three task settings, each enabled by varying the available control signals: (1) Crash Reconstruction—Given an initial image, full bounding box sequence, and a crash type, the model reconstructs a consistent video combining the visual context of the initial frame with agent motion derived from the bounding boxes. (2) Crash Prediction—Given the initial frame and only a few initial bounding box frames (e.g., 0–9), the model predicts the future motion of agents in a way that aligns with the target crash type. (3) Crash Counterfactuals—Extending the prediction task, this mode varies the crash type signal while keeping other inputs fixed, enabling the generation of multiple plausible outcomes for the same scene—supporting counterfactual safety reasoning.
@@ 3.2 Data Preparation
The data processing and preparation of crash data presented here is an essential element of our approach, allowing us to create control structures from a diverse selection of naturally occurring crashes involving cars having only a dashboard camera. Video Processing. For training, we use the MM-AU dataset (fang2024abductive, ), a large-scale collection of dashcam crash videos collected from online sources. To ensure high quality we curate this dataset by applying a series of filtering steps. We remove low-resolution or blocky videos using FFT-based heuristics (see Appendix B), detecting and excluding shot changes with PySceneDetect (PySceneDetect, ), and standardizing clips to 25-frame segments at 6 fps and $512\times 320$ resolution. Additionally, we exclude scenes involving visible humans to avoid generating violent content. For more details on the precise processing steps, please refer to Appendix C. After the filtering steps, we retain approximately 7,500 videos from the original 11,727 videos. We split these videos into a training set and a held-out test set by randomly sampling by accident type categories with a 90/10 ratio. See Appendix H for information on the dataset.
Bounding Box Extraction Pipeline. To obtain reliable bounding box annotations for all road users, we design a hybrid pipeline that combines detection and segmentation models. For detection, we use YOLOv8 (yolo2023, ) for frame-by-frame object detection. YOLO provides class-specific bounding boxes with high confidence and speed. For tracking, we use SAM2 (ravi2024sam2, ) to produce instance-level masks and reliable tracking particularly when objects get occluded or deformed, common in crash videos. This combined approach yields temporally aligned bounding boxes across all video frames. Importantly, it supports agents that enter or exit the scene dynamically—critical for realistic dynamic driving scenarios. A full breakdown of the bounding-box annotation pipeline is provided in Appendix D. We collect bounding-box annotations for all road users for the MM-AU dataset and the RUSSIA Car Crash Dataset (sivoha2023carcrash, ). These annotations along with the processed video frames are made publicly available on our project page.

### v2

@@ Abstract
Project page: https://anthonygosselin.github.io/Ctrl-Crash-ProjectPage/
keywords:
Video diffusion , Controllable generation , Autonomous Driving , Rare Event Simulation , Counterfactual Reasoning
journal: Pattern Recognition Letters
[…]
[1]organization=Mila, city=Montreal, state=Quebec, country=Canada
[…]
[2]organization=Polytechnique Montréal, city=Montreal, state=Quebec, country=Canada
[…]
[3]organization=Université de Montréal, city=Montreal, state=Quebec, country=Canada
[…]
[4]organization=McGill University, city=Montreal, state=Quebec, country=Canada
[…]
[5]organization=CIFAR AI Chair, city=Toronto, country=Canada
[…]
[6]organization=Samsung SAIL Montréal, city=Montreal, state=Quebec, country=Canada
@@ 1 Introduction
Autonomous vehicle (AV) systems must be rigorously tested in a wide range of driving situations—including rare and dangerous edge cases such as collisions to ensure safe deployment. Much of the current progress in perception, planning, and control for AVs has been driven by large-scale datasets collected in harmless, non-crash scenarios. However, realistic video data of car crashes remains extremely scarce, making it difficult to simulate, anticipate, or learn effectively from these critical events [2].
Prior work has largely approached the challenge of crash scenario modeling in two main ways.
On one hand, physics-based rendering approaches use game engines or physics simulators to model accident dynamics, but often fall short on visual realism, require expensive rendering pipelines, and demand costly human effort for environment and asset creation [3]. On the other hand, data-driven methods, such as generative models, rely on real-world footage, which is difficult to acquire in sufficient volume due to the infrequency and ethical complexity of crash events [4]. Moreover, most generative approaches focus on normal driving behavior, avoiding the complexity and unpredictability inherent in crash dynamics, where agent interactions, rare motion patterns, and semantic context all matter deeply [5, 6].
@@
Ctrl-Crash builds on latent diffusion models [7] and classifier-free guidance [8], and we extend the latter to allow independently tunable guidance strengths for each control modality, making our system highly flexible at inference time. Our two-stage training procedure first finetunes a pretrained Stable Video Diffusion (SVD) [9] model on in-the-wild ego-view accident videos, then trains a ControlNet [10] adapter to handle conditioning in order to direct the video generation. By leveraging this data-driven framework, our model can generate controllable crash videos that are visually realistic, semantically diverse, and behaviorally plausible. We see this work as a step toward not only improving the diversity and coverage of safety-critical AV testing, but also enabling counterfactual safety reasoning: the ability to simulate alternate outcomes from identical initial conditions, and better understand the causality of crashes.
@@
    We develop a data-processing pipeline to filter driving videos and extract bounding box trajectories of road users. Using it, we release a curated extension of the MM-AU dataset [11] (with bounding box annotations), as well as held-out test sets from Car Crash Dataset RUSSIA [12] (with bounding box annotations) and BDD100k [13] (with existing boxes), along with tools to support research in crash simulation and controllable video generation. Code and dataset extensions are all made open-source.
@@ 2 Related Work
Video Diffusion Models. Diffusion models [14] have emerged as a powerful paradigm for generative modeling, particularly in the domain of image synthesis. They operate by learning to reverse a gradual noising process applied to data, generating high-fidelity samples through iterative denoising. Recent advances have extended these techniques to the video domain, where temporal consistency and spatial coherence are critical [15, 9].
Given the noise-level $t$, and task-specific conditions $c$, the diffusion loss is the following:
$\mathcal{L}=\mathbb{E}_{x_{0},t,c,\epsilon\sim\mathcal{N}(0,\mathbf{I})}\Big[\|\epsilon-\epsilon_{\theta}(x_{t},t,c))\|_{2}^{2}\Big],$ (1)
where $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, $\epsilon\sim\mathcal{N}(0,\mathbf{I})$, $x_{0}$ is a training sample video, $\bar{\alpha}_{t}$ at $t\in[1,T]$ controls the diffusion schedule, and $\epsilon_{\theta}$ is a function approximator intended to predict the noise $\epsilon$ from the noise-corrupted video $x_{t}$.
Latent video diffusion models (LVDMs) [16, 9] address the computational challenges of high-resolution video generation by operating in a compressed latent space. This enables the generation of long, high-quality video sequences from compact representations. Specifically, Stable Video Diffusion (SVD) [9], an LVDM variant, leverages a UNet-based denoiser trained on video latents conditioned on an initial frame, making it suitable for tasks like image-to-video generation and temporal extension.
@@
Controllable Generative Models. Recent progress in generative modeling has emphasized not only fidelity, but also controllability — the ability to guide outputs through structured input signals. In image generation, this includes control via text prompts, sketches, bounding boxes, semantic maps, and more. In diffusion models, classifier-free guidance (CFG) [8] and ControlNet [10] have been introduced as effective methods to allow adaptable conditioning while preserving high-quality generation.
CFG is a widely-used technique in diffusion models to improve conditional generation by combining conditional and unconditional predictions, scaled to control how strongly the model follows the conditioning input. It involves jointly training the diffusion model for conditional and unconditional denoising by randomly setting the conditioning to a null value $c=\emptyset$ during training. During inference, the denoising prediction is interpolated between a conditional noise estimate $\epsilon_{\theta}(\mathbf{x}_{t},c)$ and an unconditional noise estimate $\epsilon_{\theta}(\mathbf{x}_{t},\emptyset)$, scaled by a guidance factor $\gamma$ to obtain the modified score estimate:
@@
Vision models, such as InstructPix2Pix [17], use textual or mixed-mode conditioning with CFG to manipulate generation intent and content mid-sampling. ControlNet [10] introduced an effective approach for injecting spatial control into diffusion models by adding a parallel, trainable network that processes conditioning signals and modulates the main denoising backbone. It works by branching off from intermediate layers in the main U-Net and processing a control input in parallel. Its outputs are then added to the original U-Net features before the denoising step, effectively steering the generation without retraining the base model.
@@
Car Crash Simulation and AV Safety. Video diffusion models offer a compelling solution for car crash simulation, as they can simulate both the visual realism and behavioral dynamics of complex driving scenes with very little user effort. Recent car video generation models [18, 6, 19] focus on structured driving video generation using control signals like bounding boxes or text prompts. These models generate temporally consistent driving scenes and support the synthesis of structured traffic interactions. However, they are typically limited to non-crash scenarios or coarse control.
Physics-based driving simulators [3] are useful tools for evaluating autonomous vehicle safety [20, 21, 22]. Due to the scarcity and ethical challenges of real-world crash data, there is growing interest in generating synthetic safety-critical scenarios [22, 23]. These simulations provide control and physical realism but often lack visual authenticity and broad applicability. In contrast, our work focuses on rare, adversarial cases that challenge perception and planning systems in real-world scenarios. To address the data challenge, we propose a method for processing car crash videos.
While most prior work focus on general driving scenes, notable examples of prior work tackling realistic car crash generation using real videos include:
DrivingGen [24], which generates crash videos from textual accident reports, AVD2 [25] and OAVD [11] which also generate videos of car crashes, but their main focus is car crash video description rather than generation. These methods are designed to stress-test AV policies and enhance safety coverage beyond what real-world datasets offer.
Our work builds on this vision by proposing a controllable video diffusion model that can generate crash outcomes directly from initial conditions and desired outcomes. Unlike physics simulators, Ctrl-Crash supports semantic control (e.g., crash type) and spatial trajectory specification (e.g., bounding boxes), enabling targeted stress testing of vision-based AV stacks. Moreover, our method supports counterfactual safety reasoning, allowing one to explore how small changes in agent behavior or intent could lead to dramatically different outcomes—a critical capability for understanding near-miss events and decision boundary failures. Importantly, our method has much higher visual and motion fidelity than prior works.
@@ 3 Our Method: Ctrl-Crash
Figure 2: Ctrl-Crash architecture: Ctrl-Crash treats Bounding Boxes ($B$s) as images. Both $B$s and images frames are put through a VAE image encoder. The crash type and $B$ embeddings are fed to ControlNet. The images embedding after adding noise ($x_{t}$), the noise level ($t$), and the ControlNet intermediate outputs ($c$) are fed to the Stable Video Diffusion (SVD) model to obtain the predicted noise $\epsilon_{\theta}(x_{t},t,c)$. CLIP embeddings [26] are computed by passing the first image for each video through a pretrained CLIP encoder. These CLIP embeddings are then given to the ControlNet and SVD models. The diffusion process is solved over multiple steps using classifier-free guidance in the latent space, and then decoded back to images using the VAE image decoder. See Section 3 for details.
@@ 3.1 Overview
We propose Ctrl-Crash (Figure 2), a controllable video diffusion framework designed to generate realistic car crash scenarios from a single initial frame, guided by both spatial and semantic control signals. Ctrl-Crash builds on Ctrl-V [6], a framework for generating videos from rendered bounding box trajectories, by extending its capabilities to crash-specific scenarios, offering richer control and greater flexibility. Specifically, we incorporate a new semantic control signal representing crash type and introduce a refined training procedure to handle partial and noisy conditioning.
Our method follows a two-stage training pipeline. In the first stage, we fine-tune the Stable Video Diffusion (SVD) model on the MM-AU [11] crash video dataset to improve its ability to synthesize dynamic and physically plausible driving and accident scenes. In the second stage, we train a ControlNet module to inject conditioning information in two forms: (1) bounding box sequences representing road user motion across time, and (2) discrete crash type labels encoding high-level semantic intent. To ensure generalization to incomplete or noisy control, we introduce a curriculum-based random masking strategy that progressively masks out parts of the control inputs during training. Masked bounding box frames are replaced by a learnable embedding that preserves scene plausibility. We further extend classifier-free guidance to support independent scaling of each control modality, enabling nuanced and flexible control at inference. Unconditional noise predictions are obtained from the pretrained base model for improved generation diversity and stability.
Ctrl-Crash supports three task settings, each enabled by varying the available control signals: (1) Crash Reconstruction: Given an initial image, full bounding box sequence, and a crash type, the model reconstructs a consistent video combining the visual context of the initial frame with agent motion derived from the bounding boxes. (2) Crash Prediction: Given the initial frame and either none or a few initial bounding box frames (e.g., 0–9), the model predicts the future motion of agents in a way that aligns with the target crash type. (3) Crash Counterfactuals: Extending the prediction task, this mode varies the crash type signal while keeping other inputs fixed, enabling the generation of multiple plausible outcomes for the same scene, supporting counterfactual safety reasoning (Figures 1 and 3). We use the term “counterfactual” here in an informal sense, not to imply a formal causal model, but rather to describe alternate plausible outcomes from the same initial conditions. For example, Ctrl-Crash can be applied to non-crash datasets such as BDD100K, allowing the generation of realistic crash outcomes from otherwise uneventful driving scenes (see A.2 for examples).
@@ 3.2 Data Preparation
Processing and preparing crash data is an essential element of our approach, enabling the construction of control signals from a diverse set of car crashes captured by dashboard cameras and sourced from online videos.
Video Processing. For training, we use the MM-AU dataset, a large-scale collection of dashcam crash videos collected from online sources. To ensure high quality we curate this dataset by applying a series of filtering steps. This includes automated detection and filtering of low resolution videos, shot change detection with PySceneDetect, and manual exclusion of scenes involving visible humans getting struck. We manually exclude scenes where visible humans are hit to avoid exposing the model to violent content and to prevent it from learning to depict human injury, reducing the risk of harmful or inappropriate generations post-deployment. Additionally, we standardize clips to 25-frame segments at 6 fps and $512\times 320$ resolution. After the filtering steps, we retain approximately 7,500 videos from the original 11,727 videos. We split these videos into a training set and a held-out test set by randomly sampling by accident type categories with a 90/10 ratio. For more details on the precise processing steps and the dataset, please refer to C and H.
Bounding Box Extraction Pipeline. To obtain reliable bounding box annotations for all road users, we design a hybrid pipeline that combines detection and segmentation models. For detection, we use YOLOv8 [27] for frame-by-frame object detection. YOLO provides class-specific bounding boxes with high confidence and speed. For tracking, we use SAM2 [28] to produce instance-level masks and reliable tracking particularly when objects get occluded or deformed, common in crash videos. This combined approach yields temporally aligned bounding boxes across all video frames. Importantly, it supports agents that enter or exit the scene dynamically—critical for realistic dynamic driving scenarios. A full breakdown of the bounding-box annotation pipeline is provided in D. We collect bounding-box annotations for all road users for the MM-AU dataset and the RUSSIA Car Crash Dataset [12]. These annotations along with the processed video frames will be made publicly available on our project page.